# Toward a Physical Theory of Intelligence

### v1

@@ 7.1.2 Metriplectic Decomposition and Minimal-Dissipation Argument
We now establish, under broad dynamical assumptions, that oscillatory dynamics reduce entropy production for a fixed level of goal-directed work and thereby increase intelligence under the proposed framework. This result generalises the geometric intuition of §7.1.1: just as the skew-symmetric component of a Jacobian generates volume-preserving rotations, the reversible component of a metriplectic system provides a low-dissipation mechanism for propagating internal structure and predictive information.
@@
$\dot{x}_{t}=J(x_{t},t)\,\nabla H(x_{t},t)-R(x_{t},t)\,\nabla\Xi(x_{t},t),$ (70)
where $J^{\top}=-J$ generates reversible motion and $R^{\top}=R\succeq 0$ generates irreversible relaxation. The scalar fields $H(x,t)$ and $\Xi(x,t)$ represent an internal conserved quantity and a dissipation potential, respectively. The standard GENERIC degeneracy conditions
$J(x,t)\,\nabla\Xi(x,t)=0,\qquad R(x,t)\,\nabla H(x,t)=0,$ (71)
ensure that reversible and dissipative dynamics act on orthogonal subspaces.
The generalized entropy-production rate is
$\dot{S}_{\mathrm{phys}}(t)=\big\langle\nabla\Xi(x_{t},t),R(x_{t},t)\,\nabla\Xi(x_{t},t)\big\rangle,$ (72)
@@
with corresponding irreversible-information rate
$\dot{I}_{\mathrm{irr}}(t)=\dot{S}_{\mathrm{phys}}(t)/\alpha.$ (73)
The goal potential $G(x,e,t)$ defines the instantaneous useful power
$\dot{W}_{\mathrm{goal}}(t):=-\,\tfrac{d}{dt}\,\mathbb{E}[G(x_{t},e_{t},t)],$ (74)
which decomposes under the metriplectic flow as
$\dot{W}_{\mathrm{goal}}=-\,\mathbb{E}[\nabla G^{\top}J\nabla H]+\mathbb{E}[\nabla G^{\top}R\nabla\Xi]-\mathbb{E}[\partial_{t}G].$ (75)
For a sustained task with long-run mean useful power $\overline{P}_{\mathrm{goal}}>0$, we consider recurrent dynamics that satisfy this requirement.
@@
Proposition 7.1 (Minimal irreversible information among admissible metriplectic dynamics for fixed useful work).
@@
Among all admissible metriplectic dynamics achieving a prescribed long-run mean rate of goal-directed work $\overline{\dot{W}}_{\mathrm{goal}}>0$, the long-run irreversible information rate is minimised when dissipative computation is confined to task-critical interventions and reversible dynamics are used to preserve and reuse internal structure.
@@
Time-averaging the power decomposition shows that entropy production depends only on the dissipative contribution $R\nabla\Xi$. Minimising the long-run entropy-production rate subject to a fixed useful-work constraint therefore favours dynamics in which irreversible contributions are reduced wherever possible and reversible propagation of internal structure carries the bulk of the computation. Any increase in dissipative flow beyond that required for control or intervention strictly increases the irreversible-information rate. ∎
Thus, the metriplectic decomposition partitions computation into reversible (structure-preserving) and irreversible (intervention) components. For a fixed rate of goal-directed work,
$\chi(t)=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (76)
increasing the proportion of computation carried by reversible dynamics reduces $\dot{I}_{\mathrm{irr}}(t)$ and increases intelligence. While real intelligent systems necessarily exhibit finite dissipation, the monotonic relation remains: greater oscillatory coordination reduces unnecessary dissipative loss and increases intelligence.
This principle appears across domains. Locomotion recycles mechanical conserved quantities through limit cycles, neural populations sustain information flow through phase-synchronised rhythms, and resonant electrical or mechanical systems transfer conserved quantities with minimal loss. Oscillatory dynamics therefore emerge as a dynamically privileged mechanism for reducing irreversible information loss while maintaining effective goal-directed behaviour.
7.1.3 Precision-Dissipation Tradeoff and the TUR Analogy
The minimum-dissipation principle derived above has a natural analogue in the stochastic setting. When noise is present, physical performance must be evaluated not only in terms of mean useful work but also with respect to fluctuations around that mean. The thermodynamic uncertainty relation (TUR) [3, 13] captures this tradeoff: it states that the precision of any sustained current is fundamentally limited by total entropy production in the relevant conserved-quantity channel.
$\frac{\mathrm{Var}(J_{T})}{\mathbb{E}[J_{T}]^{2}}\;\geq\;\frac{2\,\alpha}{\Sigma_{T}},\qquad\Sigma_{T}=\int_{0}^{T}\dot{S}_{\mathrm{phys}}(t)\,dt=\alpha\!\int_{0}^{T}\!\dot{I}_{\mathrm{irr}}(t)\,dt,$ (77)
where $\Sigma_{T}$ is the total generalized entropy produced. The inequality expresses a simple physical message: achieving lower relative variance requires greater dissipation. Here the TUR is used as a structural constraint illustrating the relationship between time-asymmetry, precision, and dissipation, rather than as a claim that biological systems exactly saturate a specific bound.
Although the TUR is rigorously derived for Markov processes obeying local detailed balance, its qualitative form is far more general because it ultimately reflects the same time-asymmetry that underlies irreversible information flow. Whenever the statistics of forward and time-reversed trajectories become similar, the KL divergence $D_{\mathrm{KL}}[p(x_{0:T})\|\tilde{p}(x_{0:T})]$ becomes small and the bound approaches equality.
Corollary 7.1 (Oscillatory Dynamics Approach the TUR Limit).
Internal dynamics dominated by reversible flow fields ($\dot{x}\approx J\nabla H$) are locally time-symmetric and therefore operate near equality in the TUR. For fixed mean current $\mathbb{E}[J_{T}]$, increasing the reversible (oscillatory) fraction of the dynamics reduces the generalized entropy required to maintain a given level of precision.
This stochastic perspective reinforces the deterministic minimum-dissipation argument. Oscillatory coordination suppresses time-asymmetry in microscopic trajectories, thereby minimizing the dissipation required to maintain precise, goal-directed currents in noisy environments. In this sense, reversible or near-reversible dynamics provide a particularly entropically efficient substrate for precision, complementing their role as the most efficient means of sustaining useful work in the deterministic setting.
A third line of reasoning links the GENERIC framework §2.4 to a general information-fidelity tradeoff, showing that reversible, oscillatory dynamics reduce the irreversible information rate required to sustain accurate internal representations. This argument does not rely on classical rate-distortion assumptions (e.g. stationarity or memoryless encoders); it requires only the physical fact that any system maintaining a representation of an external signal must supply a minimal flow of information to achieve a desired fidelity.
Let the system track an external process $y_{t}$ with an internal estimate $\hat{y}_{t}$ and distortion measure
$D=\mathbb{E}\!\left[d(y_{t},\hat{y}_{t})\right],$ (78)
where $d$ is any task-relevant loss function. Define $\mathcal{R}(D)$ as the minimal physically required information rate that any substrate must supply to maintain fidelity $D$. This quantity is a generalised rate-distortion function: a lower bound on the information flow needed to sustain a representation of given accuracy, independent of the mechanism by which the representation is encoded.
Within the metriplectic dynamics
$\dot{x}=J\nabla H-R\nabla\Xi,$ (79)
the reversible flow $J\nabla H$ transports predictive structure through time without generalized entropy production. This component may therefore satisfy part of the fidelity requirement $\mathcal{R}(D)$ without incurring irreversible information loss. Let $\gamma_{\mathrm{rev}}(t)$ denote the instantaneous rate at which the reversible subspace preserves predictive information, satisfying
$0\;\leq\;\gamma_{\mathrm{rev}}(t)\;\leq\;\mathcal{R}(D).$ (80)
Only the remaining portion of $\mathcal{R}(D)$ must be supplied by dissipative, irreversible updates. This yields the general bound
$\dot{I}_{\mathrm{irr}}(t)\;\geq\;\mathcal{R}(D)\;-\;\gamma_{\mathrm{rev}}(t),$ (81)
expressing that the irreversible-information rate cannot fall below the portion of the fidelity requirement not already carried by reversible predictive memory. As the reversible flow preserves more task-relevant structure, the required irreversible rate decreases toward its minimal value consistent with the task and environmental constraints.
Oscillatory dynamics, by recirculating predictive structure through phase-coherent cycles, reduce the irreversible information rate required to maintain a given representational fidelity. They therefore increase the instantaneous intelligence
$\chi(t)=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (82)
by decreasing the dissipation required for a fixed level of goal-directed performance.
In this view, the reversible subspace $J\nabla H$ functions as an internal predictive memory, transporting structure across time with minimal loss, while the dissipative flow $R\nabla\Xi$ performs corrective updates only when predictions fail. Increasing the fraction of task-relevant information preserved by reversible oscillations reduces the required dissipative updates and lowers $\dot{S}_{\mathrm{phys}}(t)$, thereby increasing intelligence.
Thus, the information-theoretic argument aligns with the energetic and stochastic perspectives developed above. Across all three interpretations—(i) minimal dissipation for fixed work, (ii) precision constraints under noise, and (iii) information-fidelity tradeoffs—the same principle emerges: oscillatory coordination enables adaptive performance to be sustained with reduced irreversible information processing. Oscillatory or phase-synchronous regimes therefore represent a dynamically privileged and informationally efficient operating regime for resource-limited intelligent systems.
7.2 Near-Critical Dynamics
Oscillatory coordination explains how biological systems reduce the irreversible information required to sustain useful work. We now examine a complementary phenomenon observed at the mesoscale: many biological systems, and neural systems in particular, operate near dynamical criticality. Empirically, neural recordings exhibit signatures of critical behaviour including long-range temporal correlations, $1/f$ spectral structure, and power-law distributed activity fluctuations [4, 26]. These observations indicate that the brain often operates in a regime poised between order and disorder rather than in a deeply subcritical or chaotic state.
Within our framework, the relevance of criticality arises from the balance it strikes between sensitivity and dissipation. At the mesoscale, an intelligent system is not evaluated by individual trajectories but by the responsiveness of its coarse-grained state distribution to small perturbations. This responsiveness can be summarized by an information-susceptibility measure quantifying how much predictive information can be acquired per unit perturbation.
Let $\dot{I}_{\mathrm{acq}}(t)$ denote the rate at which new predictive information is incorporated into the internal dynamics, and let $\dot{I}_{\mathrm{irr}}(t)$ denote the corresponding irreversible information rate. Their ratio
$\Lambda(t):=\frac{\dot{I}_{\mathrm{acq}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (83)
characterizes how much usable information is gained per unit irreversible cost. Near a critical point, coarse-grained susceptibility is high: small perturbations induce disproportionately large changes in behaviourally relevant statistics. At the same time, near-critical systems remain marginally stable, preventing runaway dissipation and keeping $\dot{I}_{\mathrm{irr}}(t)$ finite.
Proposition 7.2 (Near-critical susceptibility enhances information efficiency).
For dynamics with finite $\dot{I}_{\mathrm{irr}}(t)$, the ratio $\Lambda(t)$ is increased in regimes of high coarse-grained susceptibility. Such regimes correspond to empirical signatures of near-criticality, where information-bearing fluctuations are amplified without proportional increases in irreversible dissipation.
Proof sketch.
Increasing coarse-grained susceptibility increases the rate at which new predictive information is absorbed from small inputs, thereby increasing $\dot{I}_{\mathrm{acq}}$. Provided the dynamics remain marginally stable, the associated increase in $\dot{I}_{\mathrm{irr}}$ is comparatively smaller. The ratio $\Lambda(t)$ therefore attains elevated values in regions of maximal responsiveness under stability constraints. ∎
Near-critical systems thus provide an informational complement to the efficiency gains achieved through oscillatory coordination. Whereas oscillatory dynamics reduce the irreversible information required to sustain a fixed level of useful work, near-critical dynamics enhance the amount of behaviourally relevant information acquired per unit irreversible cost. Both mechanisms express the same overarching principle: extract the greatest possible task-relevant information while minimizing unnecessary irreversible dissipation.
@@ 7.2 Near-Critical Dynamics
In biological systems these regimes coexist and complement each other. Long-range oscillatory synchrony supports low-dissipation information propagation, while near-critical fluctuations supply heightened responsiveness and adaptive flexibility. Together, they place neural systems in an informationally efficient operating regime appropriate for intelligent, resource-limited physical agents.
@@ 8.2.3 Ports, Interfaces, and Composition Constraints
(A) 1-D Oscillator With Port Decomposition
$\dot{\theta}_{i}=\omega_{i}+v_{i}+F(\theta_{i})$ (92)
$\begin{aligned} \text{input port:}\;&v_{i}\\ \text{output port:}\;&\theta_{i}\\ \text{encoding port:}\;&\theta_{i}\\ \text{context port:}\;&\omega_{i}\end{aligned}$ (93)
(B) Coupled Oscillator Circuit
Oscillator 1Oscillator 2coupling $v_{2}$coupling $v_{1}$$\omega_{1}$$\omega_{2}$$\theta_{1}$$\theta_{2}$
(C) Coupled Oscillator State Space
Figure 2: Three views of a coupled-oscillator dynamical circuit. (A) A one-dimensional phase oscillator with input, output, logical, and context ports. (B) Bidirectional coupling of two such oscillators. (C) The resulting joint state space on the torus $S^{1}\!\times S^{1}$, showing phase-locked and drifting regimes that yield discrete logical states under CCE-compliant encoding.
@@ 8.4 Beyond Boolean Logic
Computation is fundamentally physical: digital, analog, and dynamical systems alike evolve under continuous physical laws. Classical digital architectures restrict these laws by engineering trajectories that rapidly collapse onto a small set of fixed-point attractors, with clocked gating providing transitions between equilibria. Under CCE, this corresponds to a small encoding space realised by a limited number of metastable basins and frequent irreversible transitions between them.
@@ 8.4.1 Dynamical Primitives Beyond Fixed-Point Logic
Physical dynamical systems admit invariant structures that induce qualitatively different encoding geometries under CCE, leading to distinct computational regimes and irreversible-cost profiles.
Multistability.
Systems with multiple stable equilibria induce discrete encoding sets under CCE, with computation proceeding through basin selection. Each input drives the system toward one of finitely many metastable attractors, and encoding selection occurs via irreversible collapse into a chosen basin. This regime closely resembles Boolean decision logic, but arises here as a consequence of attractor geometry rather than symbolic manipulation.
@@
Oscillatory systems encode information in phase, frequency, or amplitude through stable periodic orbits. Under CCE, phase-coherent evolution enables the reversible transport of predictive structure along limit cycles, allowing information to be preserved across time without repeated irreversible updates. For tasks with periodic or cyclic structure, this geometry substantially reduces the number of irreversible encoding transitions required to sustain accurate computation.
@@
Weakly coupled oscillators generate invariant tori, inducing continuous encoding manifolds of higher dimensionality. Such geometries support smooth interpolation between encodings and permit gradual deformation of internal representations without discrete collapse. Relative to fixed-point logic, toroidal dynamics therefore reduce irreversible merges and enable richer continuous computation under CCE.
@@
Trajectories through ordered sequences of saddle neighbourhoods realise temporally structured computation without global clocking. Under CCE, encodings unfold along directed paths in state space, with each metastable region constraining subsequent evolution. Computation is therefore carried by the ordering of transitions rather than by instantaneous state, enabling sequential processing with minimal irreversible intervention.
Hamiltonian and momentum-driven flows.
Systems governed by conserved quantities support reversible, geometry-preserving transformations generated by Hamiltonian dynamics. Under CCE, such flows transport encodings along constant-energy manifolds without entropy production, preserving information exactly in the ideal limit. Irreversibility is confined to sparse control, coupling, or readout events, making these dynamics maximally efficient substrates for information processing when task structure aligns with the conserved geometry.
Across these regimes, the central determinant of computational efficiency is the geometry of the encoding space induced by the dynamics. When this geometry aligns with task-relevant structure in the environment, useful computation can be achieved with fewer irreversible encoding transitions. Among these regimes, oscillatory dynamics provide a particularly clear and tractable example. We therefore examine frequency discrimination as a concrete case in which geometric congruence yields a provable reduction in irreversible information processing.

### v2

@@ 9.1.1 Metriplectic Decomposition and Minimal-Dissipation Argument
We now establish that oscillatory dynamics reduce generalised dissipation for a fixed level of goal-directed work, thereby strictly increasing the intelligence metric ($\chi$).
@@
$\dot{x}_{t}=J(x_{t},t)\nabla H(x_{t},t)-\mathcal{R}(x_{t},t)\nabla\Xi(x_{t},t),$ (119)
where $J^{\top}=-J$ generates reversible motion and $\mathcal{R}^{\top}=\mathcal{R}\geq 0$ generates irreversible relaxation. As established in §7, the generalised dissipation rate is
$\dot{S}_{\mathrm{phys}}(t)=\langle\nabla\Xi(x_{t},t),\mathcal{R}(x_{t},t)\nabla\Xi(x_{t},t)\rangle$ (120)
@@
with the corresponding irreversible-information rate $\dot{I}_{\mathrm{irr}}(t)=\dot{S}_{\mathrm{phys}}(t)/\alpha$.
@@
Proposition 9.1 (Minimal irreversible information among robust metriplectic dynamics).
@@
Among all structurally stable metriplectic dynamics achieving a prescribed long-run mean rate of goal-directed work $\dot{W}_{goal}>0$, the long-run irreversible information rate is minimised when dissipative computation is strictly confined to (i) maintaining the low-dimensional task manifolds (transverse amplitude contraction) and (ii) executing task-critical CCE collapses. Reversible phase dynamics ($J\nabla H$) are used to preserve and propagate internal structure along the manifolds.
Proof.
Time-averaging the power decomposition shows that generalised dissipation depends strictly on the irreversible contribution $\mathcal{R}\nabla\Xi$. For structurally stable memory transport in noisy environments, $\mathcal{R}\nabla\Xi$ must be non-zero to contract transverse noise. Minimising the long-run dissipation rate subject to robustness and useful-work constraints therefore favours geometric flows (limit cycles) that orthogonally decouple the stabilising dissipative flow from the information-carrying reversible flow. Any increase in longitudinal dissipative flow beyond the minimum required for discrete action selection strictly inflates $\dot{I}_{\mathrm{irr}}$ and lowers intelligence. ∎
9.1.2 Precision-Dissipation Tradeoff and the TUR Analogy
This minimum-dissipation principle has a natural analogue in the stochastic setting. When noise is present, physical performance must be evaluated with respect to fluctuations. The Thermodynamic Uncertainty Relation (TUR) [2, 11] captures this tradeoff, stating that the precision of any sustained current is fundamentally limited by total entropy production.
$\frac{\mathrm{Var}(J_{T})}{\mathbb{E}[J_{T}]^{2}}\geq\frac{2\alpha}{\Sigma_{T}},$ (121)
where $\Sigma_{T}$ is the total cumulative generalised dissipation over the horizon:
$\Sigma_{T}=\int_{0}^{T}\dot{S}_{\mathrm{phys}}(t)dt=\alpha\int_{0}^{T}\dot{I}_{\mathrm{irr}}(t)dt=\alpha I_{\mathrm{irr},T}.$ (122)
The inequality expresses a strict physical bound: achieving lower relative variance requires greater cumulative dissipation.
Corollary 9.1 (Oscillatory Dynamics Approach the TUR Limit).
Internal dynamics dominated by reversible flow fields $(\dot{x}\approx J\nabla H)$ are locally time-symmetric and therefore operate near equality in the TUR. For fixed mean current, increasing the reversible (oscillatory) fraction of the dynamics reduces the generalised dissipation required to maintain a given level of precision.
Oscillatory coordination suppresses time-asymmetry in microscopic trajectories, thereby minimising the physical cost required to maintain precise, goal-directed CCEs in noisy environments.
Finally, reversible, oscillatory dynamics reduce the irreversible information rate required to sustain accurate internal representations. Let the system track an external process $y_{t}$ with an internal estimate $\hat{y}_{t}$ and distortion measure $D=\mathbb{E}[d(y_{t},\hat{y}_{t})]$. Define $\mathcal{R}(D)$ as the minimal physically required information rate to maintain fidelity $D$.
Within the metriplectic dynamics, the reversible flow $J\nabla H$ transports predictive CCEs through time without generalised dissipation. Let $\gamma_{\mathrm{rev}}(t)$ denote the instantaneous rate at which the reversible subspace preserves this predictive information, satisfying $0\leq\gamma_{\mathrm{rev}}(t)\leq\mathcal{R}(D)$. Only the remaining portion must be supplied by dissipative, irreversible CCE updates. This yields the general bound:
$\dot{I}_{\mathrm{irr}}(t)\geq\mathcal{R}(D)-\gamma_{\mathrm{rev}}(t).$ (123)
Oscillatory dynamics, by recirculating predictive CCEs through phase-coherent cycles, satisfy a large portion of the fidelity requirement $\mathcal{R}(D)$ via $\gamma_{\mathrm{rev}}(t)$. This explicitly reduces the required rate of irreversible updates ($\dot{I}_{\mathrm{irr}}$), thereby increasing the instantaneous intelligence $\chi(t)$.
9.2 Near-Critical Dynamics and CCE Susceptibility
Oscillatory coordination explains how biological systems transport CCEs efficiently. We now examine near-criticality, which governs how these systems form new encodings. Neural recordings exhibit signatures of criticality, suggesting the brain operates poised between order and disorder [3, 23].
Within our framework, the relevance of criticality arises from the geometric tuning of CCE basins. To absorb new predictive information, the brain must collapse existing encodings and form new ones, incurring an irreversible cost. The efficiency of this process can be summarised by a susceptibility measure quantifying how much predictive information can be acquired per unit of generalised dissipation.
Let $\dot{I}_{\mathrm{acq}}(t)$ denote the rate at which new predictive information is incorporated, and $\dot{I}_{\mathrm{irr}}(t)$ the corresponding irreversible physical cost. Their ratio
$\Lambda(t):=\frac{\dot{I}_{\mathrm{acq}}(t)}{\dot{I}_{\mathrm{irr}}(t)}$ (124)
characterises how much usable information is gained per unit of generalised Landauer cost.
Proposition 9.2 (Near-critical susceptibility enhances encoding efficiency).
For dynamics with finite $\dot{I}_{\mathrm{irr}}(t)$, the ratio $\Lambda(t)$ is locally maximised in regimes of high coarse-grained susceptibility. This corresponds to near-criticality, where the energetic barriers between CCE basins are minimised, allowing information-bearing fluctuations to induce macroscopic encoding updates without runaway dissipation.
@@ 9.2 Near-Critical Dynamics and CCE Susceptibility
Near a critical phase transition, the energetic barriers separating metastable CCE basins become shallow. This high susceptibility allows small environmental perturbations to rapidly drive the system into new encoding states, maximising $\dot{I}_{\mathrm{acq}}$. Because the barriers are shallow, the physical generalised dissipation ($\dot{I}_{\mathrm{irr}}$) required to force this transition is minimised. Therefore, $\Lambda(t)$ attains elevated values in regions of maximal responsiveness under stability constraints. ∎
@@
In biological systems, these regimes heavily complement each other. Long-range oscillatory synchrony supports low-dissipation CCE propagation, while near-critical fluctuations supply heightened responsiveness to carve out new CCEs cheaply. Together, they place neural systems in an optimal operating regime.
@@ 10.4 Beyond Boolean Logic
Computation is fundamentally physical: digital and analog systems alike evolve under continuous physical laws. Classical digital architectures restrict these laws by engineering trajectories that rapidly collapse onto a small set of fixed-point attractors, with clocked gating providing transitions between equilibria. Under CCE, this corresponds to a small encoding space realised by a limited number of metastable basins and frequent irreversible transitions between them.
@@ 10.4.1 Dynamical Primitives Beyond Fixed-Point Logic
Physical dynamical systems admit invariant structures that induce qualitatively different encoding geometries under CCE. By altering the balance between reversible ($J\nabla H$) and dissipative ($\mathcal{R}\nabla\Xi$) flows, these distinct topologies lead to different computational regimes and irreversible-cost profiles.
Multistability (dissipation-dominated).
Systems with multiple stable equilibria induce discrete encoding sets under CCE, with computation proceeding through basin selection. Each input drives the system toward one of finitely many metastable attractors, and encoding selection occurs via the irreversible, contractive flow of $\mathcal{R}\nabla\Xi$ into a chosen basin.
This regime closely resembles classical Boolean decision logic, but arises here purely as a consequence of attractor geometry rather than symbolic manipulation. Because every transition requires full relaxation, it is thermodynamically costly.
@@
Oscillatory systems encode information in phase, frequency, or amplitude through stable periodic orbits. Under CCE, phase-coherent evolution enables the reversible transport of predictive structure along limit cycles via $J\nabla H$, allowing information to be preserved across time without repeated irreversible updates.
For tasks with periodic or cyclic structure, this geometry radically reduces the number of irreversible encoding transitions ($\dot{I}_{\mathrm{irr}}$) required to sustain accurate computation.
@@
Weakly coupled oscillators generate invariant tori, inducing continuous encoding manifolds of higher dimensionality. Such geometries support smooth interpolation between encodings and permit gradual deformation of internal representations without discrete collapse.
Relative to fixed-point logic, toroidal dynamics reduce irreversible merges and enable richer continuous computation under CCE, effectively computing over continuous variables without the need for discrete digitization.
@@
Trajectories through ordered sequences of saddle neighbourhoods realise temporally structured computation without a global clock. Under CCE, encodings unfold along directed paths in state space, with each metastable region temporarily capturing the state before instability drives it to the next.
Computation is therefore carried by the deterministic geometric ordering of transitions rather than by instantaneous state, enabling sequential processing and finite-state machine equivalents with minimal irreversible intervention.
Hamiltonian and momentum-driven flows (reversible-dominated).
Systems governed by conserved quantities support reversible, geometry-preserving transformations generated entirely by the Hamiltonian flow $J\nabla H$. Under CCE, such flows transport encodings along constant-energy manifolds without generalised dissipation, preserving information exactly in the ideal limit. Irreversibility is confined strictly to sparse control, boundary coupling, or readout events, making these dynamics maximally efficient substrates for information processing when task structure perfectly aligns with the conserved geometry.